3) Seven less than 4 times a number gives 13.
Answer #1

Let the number be x

4 imes the number is 4x

7 less than 4x is 4x-7

so 4x-7=13

4x-7=13  

4x=13+7

4x=20

x=20/4

x=5

the number is 5